A meeting of consuls was attended by 4 Americans and 2 Germans. If three consuls were selected at random, construct the probability distribution at the random variable G representing the number of Germans. Construct the probability histogram.


1
Expert's answer
2022-01-20T07:29:54-0500

Americans A = 4

Germans G = 2

Total = 6

3 out of 6 can be selected in ⁶C₃ = 20 Ways

Germans can be 0 , 1 or 2

three consuls were selected

G = 0  => ⁴C₃ x ²G₀ = 4 Ways

G = 1  => ⁴C₂ x ²G₁ = 12 Ways

G = 2  => ⁴C₁ x ²G₂ = 4 Ways 

G    P (G)           G. P(G)

0    4/20 =0.2        0

1     12/20 =0.6       0.6

2    4/20 =0.2        0.4